Jorge Cabrera paid $980 for a 15-year bond 10 years ago. The bond pays a coupon of 10 percent semiannually. Today, the bond is priced at $1,054.36. If he sold the bond today, what would be his realized yield? (Round to the nearest percent.) ** No computer answers** HP financial calculator is fine Formula is a must
Answer #1

use financial calculator as below

PV=-980 as the purchase value of the bond

FV=1054.36 is the current price or selling price

N=10*2=20 semi annual periods

PMT=1000*10%/2=50

Click CPT

Click I/Y=5.3238% per semi annual

what would be his realized yield=5.3238%*2=10.65%

the above is answer..